A bottle contains 750ml of fruit punch with a concentration of 50% pure fruit juice. Jin drinks 100ml of the punch and then refills the bottle win an equal amount of a cheaper brand of punch. If the concentration of juice in the bottle is now reduced to 48%, what was the concentration in the punch that Jill added?
Answer by ewatrrr(24785)


Hi,
*Note: the cheaper brand was added to 650ml of the 50% fruit punch
Let x be the concentration in the punch added
.50*650 + x*100 = .48*750
Simplify and solve
100x = .48*750ml - .50*650ml
100x = 360 -325
100x = 35
x = .35
Concentration in the punch that Jill added was 35%
check your answer
325 + 35 = 360